LISTSERV mailing list manager LISTSERV 16.5

Help for HEANET-FAULTS Archives


HEANET-FAULTS Archives

HEANET-FAULTS Archives


HEANET-FAULTS@LISTSERV.HEANET.IE


View:

Message:

[

First

|

Previous

|

Next

|

Last

]

By Topic:

[

First

|

Previous

|

Next

|

Last

]

By Author:

[

First

|

Previous

|

Next

|

Last

]

Font:

Proportional Font

LISTSERV Archives

LISTSERV Archives

HEANET-FAULTS Home

HEANET-FAULTS Home

HEANET-FAULTS  June 2010

HEANET-FAULTS June 2010

Subject:

HEA-NOC/20100420-2 [UPDATE] NUIG Planned Power Outage

From:

Daniel Lete <[log in to unmask]>

Reply-To:

Daniel Lete <[log in to unmask]>

Date:

Sat, 12 Jun 2010 00:21:37 +0100

Content-Type:

text/plain

Parts/Attachments:

Parts/Attachments

text/plain (122 lines)

Ticket Number: HEA-NOC/20100420-2      Ticket Status: UPDATE
Ticket Type:   scheduled               Resolver:      NUIG
Ticket Opened: 20100420 09:56 UTC+1    Problem Start: 20100611 22:00 UTC+1
Ticket Update: 20100612 00:21 UTC+1    Problem End:   20100613 12:00 UTC+1

Site/line:     CARRAROE, CASTLEBAR, DERI, GTI, IT Athlone, IT Sligo, LYIT,
               LYIT-Killybegs, Marine Institute, MARTINRYAN, mhars, NIRAN,
               NUIG, NUIG Clinical Sciences, NUIG-Gweedore, PAC, Schools
               Network, Sean, St. Angela's

Problem Description:
NUIG are expanding their Data Centre. Installation of a new 160KVA UPS
requires a new module to be added to the main Data Centre power
distribution board. This board needs to be powered off in order for the
modification to be made.

Problem Effects:
UPS and Generator and therefore all power to the data centre, including
HEAnet's PoP, will not be available during the works. The work will take
place over 8 hours, from 09.00-17.00 on Saturday 12th June 2010.


Loss of resiliency
AIT
CAO
LYIT
NIRAN
TippInst Thurles


Loss of Connectivity
GMIT (and all sub-sites)
 -Cluain Mhuire
 -Letterfrack
 -Castlebar
GTI
LYIT Killybegs
Marine Inst
NUIG (and all sub-sites)
 -DERI
 -Open Learning Centre
 -Carron Field
 -Mace Head
 -Sean Scoil Shailerna
 -Aras Shorcha Ni Ghuairim
 -Aras Mhairtin Ui Chadhain
 -Martin Ryan
 -Clinical Science
 -Gweedore
PAC
Sligo IT
STACS
TippInst Clonmel

Problem History:
20100420 09:29 BMA Opening ticket to track work. Clients will be contacted
individually to ensure date of outage is acceptable.

20100511 14:48 CS Since this outage is a problem for IT Sligo and the work
cannot be easily re-scheduled, NUIG will arrange for HEAnet kit to be
powered by the mains from nearby offices during the maintenance work
thereby avoiding a potential 8-hour long outage.

However, as not all of the HEAnet kit has dual power supplies, this will
result in two short outages: the first when the power supply is switched to
the temporary supply before maintenance begins and the second when
switching back after the work is complete.

We are currently waiting to confirm outage times with IT Sligo.

20100520 15:25 BN This work has now been scheduled to commence at 22:00 on
Friday 11th June.  A short, ~3m, outage will take place just after 22:00
when replacement power will be put in place.  Full power supply will be
swapped back in before 08:00 on the 12th with another short outage.

20100608 15:29 GL Work is to commence as planned on Friday 11th June at
22:00. However, full power supply will now be swapped back on the morning
of the 13th.

20100610 18:49 GL Updating ticket to reflect new end time

20100611 20:40 DL, power loss for 3 minutes at 20:01 as result of
preparation for work tonight. HEAnet systems recovered after power
restored.

20100612 00:09 DL, at 21:50 HEAnet systems disconnected from main power
supply and connected to alternative power supply with failed result since
alternative power supply was linked to the main. NUIG/HEAnet moves
equipment to different alternative power supply but fails due to overload.
NUIG/HEAnet connect systems to a third alternative power supply and restore
power. Systems powered up again at 23:50 and connectivity restored in
circuits affected.
NUIG will be performing testing, originally scheduled during the day of
Saturday 13 June, during the night 12-13 June. Currently awaiting for
electrical engineer to arrive in NUIG, expected at 01:00 and perform
testing upon arrival. Once completed, HEAnet systems will be restored to
original power supply and remain as before the work had started.

20100612 00:20 DL, times and events
   22:00 Switch equipment to alternative power 
   22:15 UPS Cut - equipment remains up
   22:29 Main power cut
   22:30 Alternative Power Failed
   22:57 Alternative Power restored - another alternative power sourced
   23:05 Alternative alternative power failed
   23:20 Alternative alternative power restored - too much load on one
      breaker, split to two breakers.
   00:00 Alternative alternative power stable for 40 minutes.

However not confident of stability. NUIG are to carry out emergency button
test in estimated 2 hour at 02:00 instead of 8am to 9am. HEAnet engineer
will remain onsite and then restore HEAnet equipment to mains power and
return to normal.

Time to Fix:
8 hours

-- 
This ticket can be monitored at http://www.hea.net/tickets/20100420-2

HEAnet Limited. Registered in Ireland, No. 275301.

Top of Message | Previous Page | Permalink

Advanced Options


Options

Log In

Log In

Get Password

Get Password


Search Archives

Search Archives


Subscribe or Unsubscribe

Subscribe or Unsubscribe


Archives

July 2019
June 2019
May 2019
April 2019
March 2019
February 2019
January 2019
December 2018
November 2018
October 2018
September 2018
August 2018
July 2018
June 2018
May 2018
April 2018
March 2018
February 2018
January 2018
December 2017
November 2017
October 2017
September 2017
August 2017
July 2017
June 2017
May 2017
April 2017
March 2017
February 2017
January 2017
December 2016
November 2016
October 2016
September 2016
August 2016
July 2016
June 2016
May 2016
April 2016
February 2016
January 2016
December 2015
November 2015
October 2015
September 2015
August 2015
July 2015
June 2015
May 2015
April 2015
March 2015
February 2015
January 2015
December 2014
November 2014
October 2014
September 2014
August 2014
July 2014
June 2014
May 2014
April 2014
March 2014
February 2014
January 2014
December 2013
November 2013
October 2013
September 2013
August 2013
July 2013
June 2013
May 2013
April 2013
March 2013
February 2013
January 2013
December 2012
November 2012
October 2012
September 2012
August 2012
July 2012
June 2012
May 2012
April 2012
March 2012
February 2012
January 2012
December 2011
November 2011
October 2011
September 2011
August 2011
July 2011
June 2011
May 2011
April 2011
March 2011
February 2011
January 2011
December 2010
November 2010
October 2010
September 2010
August 2010
July 2010
June 2010
May 2010
April 2010
March 2010
February 2010
January 2010
December 2009
November 2009
October 2009
September 2009
August 2009
July 2009
June 2009
May 2009
April 2009
March 2009
February 2009
January 2009
December 2008
November 2008
October 2008
September 2008
August 2008
July 2008
June 2008
May 2008
April 2008
March 2008
February 2008
January 2008
December 2007
November 2007
October 2007
September 2007
August 2007
July 2007
June 2007
May 2007
April 2007
March 2007
February 2007
January 2007
December 2006
November 2006
October 2006
September 2006
August 2006
July 2006
June 2006
May 2006
April 2006
March 2006
February 2006
January 2006
December 2005
November 2005
October 2005
September 2005
August 2005
July 2005
June 2005
May 2005
April 2005
March 2005
February 2005
January 2005
December 2004
November 2004
October 2004
September 2004
August 2004
July 2004
June 2004
May 2004
April 2004
March 2004
February 2004
January 2004
December 2003
November 2003
October 2003
September 2003
August 2003
July 2003
June 2003
May 2003
April 2003
March 2003
February 2003
January 2003
December 2002
November 2002
October 2002
September 2002
August 2002
July 2002
June 2002
May 2002
April 2002
March 2002
February 2002
January 2002
December 2001
November 2001
October 2001
September 2001
August 2001
July 2001
June 2001
May 2001
April 2001
March 2001
February 2001
January 2001
December 2000
November 2000
October 2000
September 2000
August 2000
July 2000
June 2000
May 2000
April 2000
March 2000
February 2000
January 2000
December 1999
November 1999
October 1999
September 1999
August 1999
July 1999
June 1999
May 1999
April 1999
March 1999
February 1999
January 1999
December 1998
November 1998
October 1998
September 1998
August 1998
July 1998
June 1998
May 1998
April 1998
March 1998
February 1998
January 1998
December 1997
November 1997
October 1997
September 1997
August 1997
July 1997
June 1997
May 1997
April 1997
March 1997
February 1997
January 1997
December 1996
November 1996
October 1996
September 1996
August 1996
July 1996
June 1996
May 1996
April 1996
March 1996
February 1996
January 1996
December 1995
November 1995
October 1995
September 1995
August 1995
July 1995
June 1995
May 1995
April 1995
March 1995
February 1995
January 1995
December 1994
November 1994
October 1994
September 1994
August 1994
July 1994
June 1994
May 1994
April 1994
March 1994
February 1994
January 1994
December 1993
November 1993
October 1993
September 1993
August 1993
July 1993
June 1993

ATOM RSS1 RSS2



LISTSERV.HEANET.IE

Secured by F-Secure Anti-Virus CataList Email List Search Powered by the LISTSERV Email List Manager